# Latent Space Characterization of Autoencoder Variants

### v1

@@ IV Experimental Setup
Training data: We train our models on the MNIST dataset. The CAE and VAE models are trained using 60,000 clean images, while the DAE was trained on a noisy version of the dataset, with Gaussian noise (sigma = 0.05) added to the training images.
Testing data: To effectively capture the changes in structure of the underlying latent manifold for each model, we construct a comprehensive test dataset from the MNIST test set. This dataset includes multiple classes, each containing 300 images. The first class contains clean images, while the subsequent classes are progressively corrupted with Gaussian white noise, with sigma increasing in increments of 0.01.
Training loss: For CAE and DAE, we employ a custom loss function that combines the weighted sum of MSE and SSIM losses, with equal weights. For the VAE, we train it using the MSE reconstruction loss along with the Kullback–Leibler Divergence loss, where the KLD term is weighted by a parameter $\beta=0.8$.
@@ V-A Latent space structure in manifold space
The variability in ranks of the SPSD matrices observed in CAE and DAE results in a stratified manifold. Since each SPSD matrix lies on a specific SPSD manifold defined by its rank, the latent spaces of CAE and DAE span multiple product manifolds determined by the tuple $(r_{1},r_{2},r_{3})$ creating stratification of space [26]. Moving within a stratum is smooth, but transitioning between the manifolds is not so due to the dimensionality differences between the strata. Though the DAE shows a bit more structure than the CAE in the sense that the number of strata are lesser than for CAE, it still exhibits stratification due to the rank variability of its SPSD matrices.
@@
In contrast, the VAE shows consistent ranks across all noise levels for the three covariance matrices, indicating that all tensors lie on the same product manifold - product of two SPD manifolds and an SPSD manifold. This consistency results in smooth movement within the latent space from one point to another, describing the latent space of the VAE as a smooth product manifold. As outlined in [27, Example 1.34], the Cartesian product of manifolds is a smooth product manifold if each component manifold is smooth. Since both SPD and SPSD (of fixed rank) are smooth manifolds, the latent space of the VAE can be described as a smooth product manifold. Our primary contribution lies in characterizing the latent manifold as the PM of SPD/SPSD matrices, where the core distinction between smooth and non-smooth manifolds is rooted in the structure of this product manifold. CAE and DAE form stratified product manifolds due to rank variability of their respective SPSD matrices, leading to discontinuities and non-smooth transitions. On the other hand, the VAE, with its uniform rank structure, results in a smooth product manifold.
@@ V-B Latent space structure in Hilbert space
We observe that for the CAE and DAE, the dimensionality of subspaces decreases as the input transitions from clean to noisy, indicating that the subspaces in the Hilbert space change with increasing noise. The CAE experiences a sharper drop in dimensionality, while DAE preserves it slightly better. In contrast, the VAE points lie in the same subspace regardless of the noise level.
We also examine how PSNR behaves across different noise levels (Fig. 4). It is observed that as the subspace dimension decreases, PSNR tends to drop, particularly in CAE and DAE, whereas VAE maintains both constant dimensionality and consistent PSNR across all noise levels, suggesting a connection between the two.
With noisy subspace dimensions differing from those of the clean subspace, we have so far established that CAE and DAE points lie on distinct subspaces for noisy cases. To examine how the subspaces corresponding to noisy inputs are oriented with respect to the clean ones, we calculate the principal angles [28] between noisy and clean subspaces at each noise level. Given two subspaces $\mathcal{X}$ and $\mathcal{X^{\prime}}$ with dimensions $d$ and $d^{\prime}$, respectively, the number of principal angles is determined by $m=\min(d,d^{\prime})$. The results, presented in Fig.5, show that for CAE and DAE, the principal angles increase with noise level, suggesting that the noisy subspaces diverge away from the clean ones with noise. This divergence is more pronounced in the CAE. In contrast, the VAE shows zero principal angles as expected.
@@ V-C Visualizing the latent tensors
We observe that as noise increases, the latent tensors for CAE and DAE diverge further from the clean points, thus confirming our earlier observation from the Hilbert space analysis, where the principal angles between noisy and clean subspaces consistently increased with higher noise levels. In DAE, the divergence is more gradual compared to CAE. In contrast, for VAE, the t-SNE coordinate range shows that the divergence between clean and noisy points is quite small, with both clean and noisy representations remaining within a single, tightly clustered region centered around the origin. These observations reinforce the behaviour we observed in both the manifold and Hilbert space analyses.
VI Conclusion
We characterize the latent spaces of different autoencoder models, specifically CAEs, DAEs and VAEs, to gain an understanding of their respective smoothness properties. For this, we explore the latent space in two domains: the manifold space and the Hilbert space. In the manifold space, with the help of a simple observation that the latent tensors lie on a PM of the SPSD matrices, we observe the variability in ranks of the SPSD matrices in CAE and DAE that result in a stratified structure, where each stratum is smooth but the overall structure is non-smooth due to discontinuities between strata. In contrast, VAE shows consistent ranks, forming a smooth product manifold. In the Hilbert space, varying dimensions and increasing principal angles between clean and noisy subspaces in CAE and DAE suggest distinct subspaces for clean and noisy data, while VAE maintains the subspace with the same dimensionality for both, with zero principal angles. We also note a close relationship between subspace dimensionality and reconstruction performance across models. These results are corroborated by the t-SNE plots where significant divergence is observed between clean and noisy points in CAE and DAE, while VAE’s points are tightly clustered near the origin.

### v2

@@ IV Experimental Setup
Training data: We train our models on the MNIST dataset. The CAE and VAE models are trained using 60,000 clean images, while the DAE is trained on a noisy version of the dataset, with Gaussian noise (sigma = 0.05) added to the training images.
Testing data: To effectively capture the changes in structure of the underlying latent manifold for each model, we construct a comprehensive test dataset from the MNIST test set. This dataset includes multiple classes, each containing 300 images. The first class contains clean images, while the subsequent classes are progressively corrupted with Gaussian white noise, with the variance increasing in increments of 0.01.
Training loss: For the CAE and DAE, we employ a custom loss function that combines the weighted sum of MSE and SSIM losses, with equal weights. For the VAE, we train it using the MSE reconstruction loss along with the Kullback–Leibler Divergence loss, where the KLD term is weighted by a parameter $\beta=0.8$.
@@ V-A Latent space structure in manifold space
The variability in the ranks of the SPSD matrices observed in CAE and DAE results in a stratified manifold. Since each SPSD matrix lies on a specific SPSD manifold defined by its rank, the latent spaces of the CAE and DAE span multiple product manifolds determined by the tuple $(r_{1},r_{2},r_{3})$ creating stratification of space [26]. In contrast, the VAE shows consistent ranks across all noise levels for the three covariance matrices, indicating that all tensors lie on the same product manifold - product of two SPD manifolds and an SPSD manifold. This consistency results in smooth movement within the latent space from one point to another, describing the latent space of the VAE as a smooth product manifold.
@@ V-B Latent space structure in Hilbert space
We observe that for the CAE and DAE, the dimensionality of subspaces decreases as the input transitions from clean to noisy, indicating that the subspaces in the Hilbert space change with increasing noise. The CAE experiences a sharper drop in dimensionality, while the DAE preserves it slightly better. In contrast, the VAE points lie in the same subspace regardless of the noise level.
With noisy subspace dimensions differing from those of the clean subspace, we have so far established that CAE and DAE points lie on distinct subspaces for noisy cases. To examine how the subspaces corresponding to noisy inputs are oriented with respect to the clean ones, we calculate the principal angles [28] between noisy and clean subspaces at each noise level. Given two subspaces $\mathcal{X}$ and $\mathcal{X^{\prime}}$ with dimensions $d$ and $d^{\prime}$, respectively, the number of principal angles is determined by $m=\min(d,d^{\prime})$. The results, presented in Fig.5, show that for the CAE and DAE, the principal angles increase with noise level, suggesting that the noisy subspaces diverge away from the clean ones with noise. This divergence is more pronounced in the CAE. In contrast, the VAE shows zero principal angles as expected.
We also examine how PSNR behaves across different noise levels (Fig. 4). It is observed that as the subspace dimension decreases, the PSNR tends to drop, particularly in CAE and DAE, whereas VAE maintains both constant dimensionality and consistent PSNR across all noise levels, suggesting a connection between the two.
@@ V-C Visualizing the latent tensors
We observe that as noise increases, the latent tensors for CAE and DAE diverge further from the clean points, thus confirming our earlier observation from the Hilbert space analysis, where the principal angles between noisy and clean subspaces consistently increased with higher noise levels. In DAE, the divergence is more gradual compared to CAE. In contrast, for VAE, the t-SNE coordinate range shows that the divergence between clean and noisy points is quite small, with both clean and noisy representations remaining within a single, tightly clustered region centered around the origin. These observations reinforce the behaviour that we observe in both the manifold and Hilbert space analyses.
@@
V-D Discussion
In this subsection, we discuss the implications of the observations from the results discussed earlier in this section, focusing on their relevance to the denoising application from the latent space perspective.
It is observed that the CAE and DAE have a latent space which is a stratified manifold. While moving within a stratum is smooth, transitioning across strata is not so due to the dimensionality differences among the strata. It is interesting to note that with increasing noise, the manifold becomes a smooth SPD manifold and moving from this higher rank manifold of SPD matrices to the lower rank manifold of SPSD matrices is easier as it only involves thresholding of the smaller eigenvalues. This observation points to a possibility of using both the autoencoders for denoising.
For VAEs, as outlined in [27, Example 1.34], the Cartesian product of manifolds is a smooth product manifold if each component manifold is smooth. Since both SPD and SPSD (of fixed rank) are smooth manifolds, the latent space of the VAE can be described as a smooth product manifold. This smooth manifold structure may lead to far simpler denoising algorithms than those with the CAEs and DAEs. Our primary contribution lies in characterizing the latent manifold as a product manifold of the SPD/SPSD matrices, where the core distinction between smooth and non-smooth manifolds is rooted in the structure of this product manifold. The CAE and DAE form stratified product manifolds due to rank variability of their respective SPSD matrices, leading to discontinuities and non-smooth transitions across strata. On the other hand, the VAE, with its uniform rank structure, results in a smooth product manifold.
The Hilbert space analysis for the three autoencoders reveals the following structure. For the CAE and DAE, the subspace representing noisy data rotates away from that of clean data and has a lower dimension. Whereas, for the VAE, the noisy and clean data points are transformed to latent space points that can be mapped to the same Hilbert space. The t-SNE plots also reflect these distinct structural differences in latent spaces of the three autoencoders. The t-SNE plot of VAE feature-points appear as a coherent ball despite varying noise levels. Contrary to this, these points are seen as elongated points in the t-SNE plot for the CAE. The DAE, being midway between these two, has more tightly clustered points.
All the three analysis methods provide strong and concurring evidence for the smoothness of the latent space of VAE and the non-smooth structure of latent space of CAE and DAE.
VI CONCLUSION AND FUTURE WORK
We characterize the latent spaces of different autoencoder models, specifically CAEs, DAEs and VAEs, to gain an understanding of their respective smoothness properties. For this, we explore the latent space in two domains: the manifold space and the Hilbert space. In the manifold space, with the help of a simple observation that the latent tensors lie on a product manifold of the SPSD matrices, we observe the variability in ranks of the SPSD matrices in the CAE and DAE that result in a stratified structure, where each stratum is smooth but the overall structure is non-smooth due to discontinuities among strata. In contrast, the VAE shows consistent ranks, forming a smooth product manifold. In the Hilbert space, varying dimensions and increasing principal angles between clean and noisy subspaces in the CAE and DAE suggest distinct subspaces for clean and noisy data, while the VAE maintains the subspace with the same dimensionality for both, with zero principal angles. We also note a close relationship between subspace dimensionality and reconstruction performance across models. These results are corroborated by the t-SNE plots where significant divergence is observed between clean and noisy points in the CAE and DAE, while such points for the VAE are tightly clustered near the origin.
Much work remains to be done. We plan to extend this analysis to other autoencoder variants and validate the results with more datasets. Further, we also plan to characterize the latent spaces of other generative models. As a constructive next step, we intend to devise denoising and deblurring algorithms by leveraging the understanding of the manifold structure of autoencoders, particularly VAEs.